# Automated Compilation Including Dropouts: Tolerating Defective Components in Stabiliser Codes

### v1

@@ Abstract
Utility-scale solid-state quantum devices will need to fabricate quantum devices at scale using imperfect processes. By introducing tolerance to fabrication defects into the design of the quantum devices, we can improve the yield of usable quantum chips and lower the cost of useful systems. Automated Compilation Including Dropouts (ACID) is a framework that works in the ancilla-free (or ‘middle-out’) paradigm, to generate syndrome extraction circuits for general stabiliser codes in the presence of defective couplers or qubits. In the ancilla-free paradigm, we do not designate particular qubits as measurement ancillas, instead measuring stabilisers using any of the data qubits in their support. This approach leads to a great deal of flexibility in how syndrome extraction circuits can be implemented. Given a stabiliser code and information about the functional and defective couplers between the qubits, ACID works by constructing and solving an optimisation problem within the ancilla-free paradigm to find a short syndrome extraction circuit. Applied to the surface code, ACID produces syndrome-extraction circuits of depth between $1\times$ (no overhead) and $1.5\times$ relative to the depth of defect-free circuits. The LUCI algorithm, the best prior art, yielded a $2\times$ overhead, so ACID offers a significant time saving. The yield of surface code chips with a logical error rate at most $10\times$ the dropout-free baseline is up to $3\times$ higher using ACID than using LUCI. I demonstrate the broad applicability of ACID by compiling syndrome extraction circuits for bivariate bicycle codes and the colour code. For these circuits, we incur a circuit-depth overhead of between $1\times$ (no overhead) and $2.5\times$ relative to defect-free circuits. Depending on the underlying connectivity, we can often accommodate isolated defective couplers and qubits with only an order-unity impact on the logical error rate. I believe this work is the first to simulate both of these families of codes in the presence of fabrication defects.
@@ 1 Introduction
Since these quantum error-correcting codes were created to handle imperfections in circuit execution, it would seem natural that they could also handle known, permanent imperfections in the quantum devices themselves. Even state-of-the-art fabrication techniques do not have a perfect yield [aruteQuantumSupremacyUsing2019, acharyaQuantumErrorCorrection2025b], and requiring that each quantum chip is perfect—that every qubit and every coupler between the qubits is functional and high fidelity—would incur a substantial, perhaps unviable, manufacturing overhead in post-selecting for chips that turned out perfectly. I note that the classical semiconductor industry, whose products are far less delicate devices than transmons or SETs, routinely builds redundancy into its chip architectures to accommodate inevitable defects. In the same way that QECCs amplify improvements in gate execution, some level of defect-tolerance in the architecture of QPUs could substantially amplify the gain in device yield afforded by improvements in fabrication.
@@
There is plenty of literature proposing techniques to handle dropouts (as I will call them henceforth) in the surface code [augerFaulttoleranceThresholdsSurface2017, siegelAdaptiveSurfaceCode2023, strikisQuantumComputingScalable2023, linCodesignQuantumErrorcorrecting2024]. The state-of-the-art is the LUCI framework [debroyLUCISurfaceCode2024, higgottHandlingFabricationDefects2025], which can handle multiple dropped couplers and qubits, often without losing code distance. However, LUCI syndrome extraction circuits have a circuit depth twice that in the dropout-free case. This results in an overall increase in computation time by at least the same factor, and in addition substantially increases logical error rates as each round provides a longer window for errors to accrue.
@@
Applied to the surface code, ACID maintains code-capacity distance in the same situations that LUCI does, and can typically find syndrome extraction cycles that have either the same depth or $1.5\times$ the depth of the dropout-free circuits—a reduction in time-overhead of between $25\%$ and $50\%$. Its distance-preserving properties are comparable: depending on whether we assume an underlying square-grid (degree-4) or hex-grid (degree-3) connectivity, we can tolerate the loss of certain qubits and couplers without losing distance. I implemented a version of the LUCI algorithm as a special case of ACID with additional constraints (Appendix E), and find that ACID is able to maintain error rates within an order of magnitude of the dropout-free baseline up to $3\times$ more often than LUCI for small numbers of dropped qubits and couplers (Figure 1). I demonstrate ACID applied to the bivariate bicycle code assuming both the degree-5 connectivity introduced in [shawLoweringConnectivityRequirements2025], and an alternative degree-6 connectivity that introduces more redundancy and leads to greater robustness in qubit dropouts. These result in syndrome extraction circuits of depth between $1\times$ and $2.5\times$ the dropout-free case, depending on the number of dropouts. I applied ACID to both the ‘middle-out’ and ‘superdense’ formulations of the colour code presented in [gidneyNewCircuitsOpen2023a], trying the former on both a hex-grid and square-grid lattice and the latter on a square-grid lattice. Again, the syndrome extraction circuits are of length between $1\times$ and $2.5\times$ the dropout-free circuits, with the three configurations each offering a different balance of qubit overhead, syndrome extraction circuit depth, and logical error rate.
@@
The source code for ACID, with examples, is available on GitHub [RiverlaneACID]. The repository also hosts visualisations of syndrome extraction circuits produced by the algorithm.
@@
Figure 1: A comparison between ACID and the state-of-the-art LUCI algorithm [debroyLUCISurfaceCode2024] when both are applied to the distance-11 surface code. Top left: the proportion of cases where ACID and LUCI are able to produce syndrome extraction circuits with a logical error rate less than $10^{-6}$ for a distance-11 surface code implemented on a device with square-grid connectivity, assuming a random subset of qubits and couplers are defective. The numbers of defective qubits and defective couplers are shown on the $x$-axis, and the shaded region is the standard binomial error. Bottom left: In the same situations, the average number of ‘contraction layers’ (Section 2) required to construct a round of syndrome extraction, plotted with the standard error in the mean. LUCI always requires 4 rounds, whereas ACID is observed to never require more than 3. Top right and bottom right: The same comparison for the distance-11 surface code implemented on a device with hex-grid connectivity. These plots are a summary of the full data provided in Appendix D.
@@ 2 Ancilla-free QEC
Vasmer and Kubica introduced the ‘morphing’ technique for stabiliser codes, in which a subset of the physical qubits in a stabiliser code are disentangled from the rest using a unitary circuit, leaving behind a different code that contains the same logical information but has fewer physical qubits [vasmerMorphingQuantumCodes2022] and different error-correction performance. The ancilla-free (or ‘middle-out’) view of QEC was first proposed in [mcewenRelaxingHardwareRequirements2023a] for the surface code, generalised to the colour code in [gidneyNewCircuitsOpen2023a] and generalised to the bivariate bicycle codes in [shawLoweringConnectivityRequirements2025]. All three works can be seen as an application of the original morphing technique. In the ancilla-free framework, we map a Calderbank-Shor-Steane (CSS) stabiliser code with $n$ qubits onto $n$ physical qubits, without any additional qubits designated for measuring the stabilisers. In other words, stabilisers are defined as subsets of the physical qubits along with a Pauli type (XX/$Z$).
The ancilla-free framework in general requires the same number of physical qubits as in ancilla-based QEC to implement codes with the same distance and logical qubits, but has the key advantage that it allows far more flexibility in constructing syndrome extraction circuits. In [mcewenRelaxingHardwareRequirements2023a] and [shawLoweringConnectivityRequirements2025] this flexibility is leveraged to reduce the connectivity required to implement surface and bivariate bicycle codes. In the LUCI framework [debroyLUCISurfaceCode2024, higgottHandlingFabricationDefects2025] and in the present work, this flexibility allows us to work around dropouts (and optionally maintain the reduced connectivity requirements).
@@
In the next contraction layer, we choose a different (not necessarily disjoint) set of stabilisers to contract, and repeat the process of contracting these onto some chosen root qubits, measuring and resetting the root qubits, and undoing the contraction. We perform as many of these layers as we need to measure each stabiliser at least once. Note that in some cases, the ancilla-free and ancilla-based paradigms can describe the same syndrome extraction circuit: notably, in the most basic distance-$d$ rotated surface code circuits presented in [mcewenRelaxingHardwareRequirements2023a], the mid-cycle code has $\sim 2d^{2}$ physical qubits, and each end-cycle code can be viewed as an unrotated surface code supported on $\sim d^{2}$ physical ‘data’ qubits alongside $\sim d^{2}$ disentangled ‘ancilla’ qubits. In this work, we privilege the mid-cycle code, fixing it and allowing the end-cycle codes to vary based on the connectivity and dropouts.
@@
As depicted in Figure 2b, each stabiliser can usually be contracted and measured in multiple ways, and this redundancy can be exploited to measure stabilisers despite the presence of broken couplers. There are cases (e.g. the bottom panel of Figure 2b) where the presence of dropped couplers and/or qubits causes the support of stabilisers to be reduced, or even causes stabilisers to split into multiple parts. When this happens, the objects produced may no longer commute with each other, or with the undamaged stabilisers. In this case we deploy the theory of subsystem codes [baconOperatorQuantumError2006], allowing us to define product stabilisers made up of products of the anticommuting ‘quasi-stabilisers’ that do commute with one another.
@@ 3.1 Constructing a subsystem code from a defective stabiliser code
Figure 4: The commutation relations of a CSS subsystem code (symplectic basis). The left and right ‘columns’ are each binary matrices, each of width $2n$ (twice the number of physical qubits) and height $n$, so that, taken together, the $2n$ rows form a complete basis for $F_{2}^{2n}$. Each row of either column represents a Pauli operator. Within each row the first $n$ bits represent the places where a Pauli operator contains an $X$, and the last $n$ bits represent where it contains a $Z$. Then the diagram can be read as a series of left-right pairs of Pauli operators, so that each operator anticommutes with its neighbour and commutes with all others. Within this representation, there are blocks corresponding to the parity check matrices $H^{X}$ and $H^{Z}$, the gauge operators $G^{X}$ and $G^{Z}$, and logical operators $L^{X}$ and $L^{Z}$. To complete the basis, I’ve included the destabiliser matrices $D^{X}$ and $D^{Z}$, although for practical purposes we do not need to compute these. The bulk of the stabilisers are the existing untouched stabilisers $U^{X}$ and $U^{Z}$, with the remaining stabilisers and gauge operators formed as products of the ‘touched’ quasi-stabilisers. Viewed this way, it is clear that a subsystem code can be seen as a stabiliser code with an arbitrary partition of the logical qubits into both logical qubits and gauge qubits. The other definition of a subsystem code, via a gauge group, is illustrated with a grey box. (Note - this representation ignores phases, which we would need to include if we were to write a full stabiliser tableau—see [nielsenQuantumComputationQuantum2010])
@@ 3.2 Producing contraction schedules and finding scheduling constraints
Contraction schedules can be uniquely specified by 1) a rooted-spanning-tree subgraph of the local graph, and 2) a ‘timing’: for each edge $e$, a timestep $1\leq t_{e}\leq t$, such that no two edges that share a node are contracted at the same time, and contractions over edges more distant from the root occur before those closer to the root. In the case of $X$-type quasi-stabilisers, this means that the first CNOTs to occur are targeted on leaf nodes of the tree, and remove the support of the Pauli string on the leaf-node qubits. Subsequent CNOTs then recede the support further up the tree until it is localised at the root. Note that some rooted spanning trees have no valid timings. The spanning trees can be enumerated in linear time using a standard algorithm [winterAlgorithmEnumerationSpanning1986]. The timings can also be enumerated in linear time using a simple recursive algorithm (see the source).
@@ 3.3 Constructing an integer linear programming problem
ACID solves the problem using CP-SAT [GoogleOrtools2025], which is known to handle large-scale scheduling problems well. The solver is given the task of trying to minimise the number of layers required to measure all the stabilisers at least once over a round. As a secondary heuristic optimisation, it tries to maximise the number of times each stabiliser is measured. The output of the solver gives us $L$, and a ‘global schedule’—a recipe for which quasi-stabilisers are to be measured in which layers and using which contraction schedule.
@@ 3.4 Assigning detectors
Stim [gidneyStimFastStabilizer2021b], the Clifford simulator used here for the memory experiments, defines ‘detectors’ as sets of measurements whose overall parity is deterministic when no errors occur: violations of these expected parities are reported to the decoder. It is the task of the decoder to infer from the detector results which errors occurred in the circuit.
It is most helpful to think of these detectors in terms of Pauli flows [browneGeneralizedFlowDeterminism2007]. For our purposes, Pauli flows are Pauli strings that propagate through our circuit in the usual way, where we can add to the support of the Pauli string (‘seed the flow’) and where we can reduce the support of the Pauli strings (‘terminate the flow’) on measurements in the correct basis. Any Pauli flow that is both seeded and terminated within the circuit can be used to define a valid detector.
@@ 4 The surface code
Figure 6: A LUCI-like diagram showing a syndrome extraction circuit of the brickwork surface code in the presence of two dropped qubits and two dropped couplers consisting of three contraction layers. I’ve followed the notation of [debroyLUCISurfaceCode2024], where stabilisers that are contacted in a given layer are indicated with a black depiction of the rooted spanning tree used to contract them. Dropped qubits and couplers are indicated in red. Click here to open the circuit in Shatter.
The state of the art for handling dropouts in the surface code is the LUCI framework [debroyLUCISurfaceCode2024], which was the original inspiration for this work. LUCI offers an intuitive way of measuring all quasi-stabilisers over $L=4$ contraction layers within a round of syndrome extraction. The technique presented in the original paper relied on the flexibility in contraction schedules offered by a (degree-4) square-grid connectivity to deal with dropped couplers and resorted to forming product stabilisers to deal with dropped qubits, with certain pathological patterns of dropped couplers treated as dropped qubits. The more recent work [higgottHandlingFabricationDefects2025] allows LUCI to be applied to a degree-3 hexagonal/brickwork-style layout by assigning product stabilisers as soon as couplers are dropped (as well as some scheduling improvements). In both works, couplers can often be dropped without reducing the code-capacity distance $d$ of the code. In particular, in the square-grid case the loss of an isolated coupler can always be accommodated without losing distance, and in the hex-grid case we can tolerate the loss of about a third of the couplers without losing distance, with the distance reducing by one for the other couplers.

### v2

@@ Abstract
Utility-scale solid-state quantum devices will need to fabricate quantum devices at scale using imperfect processes. By introducing tolerance to fabrication defects into the design of the quantum devices, we can improve the yield of usable quantum chips and lower the cost of useful systems. Automated Compilation Including Dropouts (ACID) is a framework that works in the ancilla-free (or ‘middle-out’) paradigm, to generate syndrome extraction circuits for general stabiliser codes in the presence of defective couplers or qubits. In the ancilla-free paradigm, we do not designate particular qubits as measurement ancillas, instead measuring stabilisers using any of the data qubits in their support. This approach leads to a great deal of flexibility in how syndrome extraction circuits can be implemented. Given a stabiliser code and information about the functional and defective couplers between the qubits, ACID works by constructing and solving an optimisation problem within the ancilla-free paradigm to find a short syndrome extraction circuit. Applied to the surface code, ACID produces syndrome extraction circuits of depth between $1\times$ (no overhead) and $1.5\times$ relative to the depth of defect-free circuits. The original LUCI algorithm, the best prior art, yielded a $2\times$ overhead, so ACID offers a significant time saving. The yield of surface code chips with a logical error rate at most $10\times$ the dropout-free baseline is up to $3\times$ higher using ACID than using original LUCI. I demonstrate the broad applicability of ACID by compiling syndrome extraction circuits for bivariate bicycle codes and the colour code. For these circuits, we incur a circuit-depth overhead of between $1\times$ (no overhead) and $2.5\times$ relative to defect-free circuits. Depending on the underlying connectivity, we can often accommodate isolated defective couplers and qubits with only an order-unity impact on the logical error rate. I believe this work is the first to simulate both of these families of codes in the presence of fabrication defects.
@@ 1 Introduction
Since these quantum error-correcting codes were created to handle imperfections in circuit execution, it would seem natural that they could also handle known, permanent imperfections in the quantum devices themselves. Even state-of-the-art fabrication techniques do not have a perfect yield [3, 1], and requiring that each quantum chip is perfect—that every qubit and every coupler between the qubits is functional and high fidelity—would incur a substantial, perhaps unviable, manufacturing overhead in post-selecting for chips that turned out perfectly. I note that the classical semiconductor industry, whose products are far less delicate devices than transmons or SETs, routinely builds redundancy into its chip architectures to accommodate inevitable defects. In the same way that QECCs amplify improvements in gate execution, some level of defect-tolerance in the architecture of QPUs could substantially amplify the gain in device yield afforded by improvements in fabrication.
@@
There is plenty of literature proposing techniques to handle dropouts (as I will call them henceforth) in the surface code [4, 25, 26, 17, 15] (see also [20], where dropouts on Floquet codes are explored). The state-of-the-art is the LUCI framework [8, 13], which can handle multiple dropped couplers and qubits, often without losing code distance. However, LUCI syndrome extraction circuits have a circuit depth twice that in the dropout-free case. This results in an overall increase in computation time by at least the same factor, and in addition may increase logical error rates as each round provides a longer window for errors to accrue.
@@
I demonstrate ACID applied to the bivariate bicycle code assuming both the degree-5 connectivity introduced in [24], and an alternative degree-6 connectivity that introduces more redundancy and leads to greater robustness in qubit dropouts. These result in syndrome extraction circuits of depth between $1\times$ and $2.5\times$ the dropout-free case, depending on the number of dropouts. I applied ACID to both the ‘middle-out’ and ‘superdense’ formulations of the colour code presented in [10], trying the former on both a hex-grid and square-grid lattice and the latter on a square-grid lattice. Again, the syndrome extraction circuits are of length between $1\times$ and $2.5\times$ the dropout-free circuits, with the three configurations each offering a different balance of qubit overhead, syndrome extraction circuit depth, and logical error rate.
Applied to the surface code, ACID maintains code-capacity distance in the same situations that LUCI does, and can typically find syndrome extraction cycles that have either the same depth or $1.5\times$ the depth of the dropout-free circuits—a reduction in time-overhead of between $25\%$ and $50\%$. Its distance-preserving properties are comparable: depending on whether we assume an underlying square-grid (degree-4) or hex-grid (degree-3) connectivity, we can tolerate the loss of certain qubits and couplers without losing distance. I implemented a version of the LUCI algorithm as a special case of ACID with additional constraints (Appendix E), and find that ACID is able to maintain error rates within an order of magnitude of the dropout-free baseline up to $3\times$ more often than LUCI for small numbers of dropped qubits and couplers (Figure 1).
@@
The source code for ACID, with examples, is available on GitHub [23]. The repository also hosts visualisations of syndrome extraction circuits produced by the algorithm.
Update (January 2026): the comparison made here is to a reproduction of LUCI insofar as it is detailed in the original two works [8, 13]. Shortly after the release of this manuscript, Anker and Debroy [2] showed by means of a direct comparison that ACID produces surface code syndrome extraction circuits with higher error rates than those produced by the closed-source implementation of LUCI upon which the original LUCI papers were based, explaining that that their implementation contains heuristic optimisations that are not presented in the original work. Anker and Debroy also detail a new version of LUCI using very similar techniques to those in the present work, but equipped with a set of additional, explicitly described optimisation heuristics, meaning the most up-to-date version of the LUCI algorithm is both more performant and more reproducible than previous versions, and is also capable of matching ACID’s time overhead. In my view, the newest version of LUCI should be considered the state-of-the-art for tolerating fabrication defects in the surface code. Promisingly, the heuristics presented therein are compatible with the general framework of ACID, and future work should merge the two with the aim of producing similarly optimised syndrome extraction circuits for other stabiliser codes.
@@
Figure 1: A comparison between ACID and the state-of-the-art LUCI algorithm [8] when both are applied to the distance-11 surface code. Top left: the proportion of cases where ACID and LUCI are able to produce syndrome extraction circuits with a logical error rate less than $10^{-6}$ for a distance-11 surface code implemented on a device with square-grid connectivity, assuming a random subset of qubits and couplers are defective. The numbers of defective qubits and defective couplers are shown on the $x$-axis, and the shaded region is the standard binomial error. Bottom left: In the same situations, the average number of ‘contraction layers’ (Section 2) required to construct a round of syndrome extraction, plotted with the standard error in the mean. LUCI always requires 4 rounds, whereas ACID is observed to never require more than 3. Top right and bottom right: The same comparison for the distance-11 surface code implemented on a device with hex-grid connectivity. These plots are a summary of the full data provided in Appendix D.
@@ 2 Ancilla-free QEC
Vasmer and Kubica introduced the ‘morphing’ technique for stabiliser codes, in which a subset of the physical qubits in a stabiliser code are disentangled from the rest using a unitary circuit, leaving behind a different code that contains the same logical information but has fewer physical qubits [27] and different error-correction performance. The ancilla-free (or ‘middle-out’) view of QEC was first proposed in [19] for the surface code, generalised to the colour code in [10] and generalised to the bivariate bicycle codes in [24]. All three works can be seen as an application of the original morphing technique. In the ancilla-free framework, we map a Calderbank-Shor-Steane (CSS) stabiliser code with $n$ qubits onto $n$ physical qubits, without any additional qubits designated for measuring the stabilisers. In other words, stabilisers are defined as subsets of the physical qubits along with a Pauli type (XX/$Z$).
The ancilla-free framework in general requires the same number of physical qubits as in ancilla-based QEC to implement codes with the same distance and logical qubits, but has the key advantage that it allows far more flexibility in constructing syndrome extraction circuits. In [19] and [24] this flexibility is leveraged to reduce the connectivity required to implement surface and bivariate bicycle codes. In the LUCI framework [8, 13] and in the present work, this flexibility allows us to work around dropouts (and optionally maintain the reduced connectivity requirements).
@@
In the next contraction layer, we choose a different (not necessarily disjoint) set of stabilisers to contract, and repeat the process of contracting these onto some chosen root qubits, measuring and resetting the root qubits, and undoing the contraction. We perform as many of these layers as we need to measure each stabiliser at least once. Note that in some cases, the ancilla-free and ancilla-based paradigms can describe the same syndrome extraction circuit: notably, in the most basic distance-$d$ rotated surface code circuits presented in [19], the mid-cycle code has $\sim 2d^{2}$ physical qubits, and each end-cycle code can be viewed as an unrotated surface code supported on $\sim d^{2}$ physical ‘data’ qubits alongside $\sim d^{2}$ disentangled ‘ancilla’ qubits. In this work, we privilege the mid-cycle code, fixing it and allowing the end-cycle codes to vary based on the connectivity and dropouts.
@@
As depicted in Figure 2b, each stabiliser can usually be contracted and measured in multiple ways, and this redundancy can be exploited to measure stabilisers despite the presence of broken couplers. There are cases (e.g. the bottom panel of Figure 2b) where the presence of dropped couplers and/or qubits causes the support of stabilisers to be reduced, or even causes stabilisers to split into multiple parts. When this happens, the objects produced may no longer commute with each other, or with the undamaged stabilisers. In this case we deploy the theory of subsystem codes [5], allowing us to define product stabilisers made up of products of the anticommuting ‘quasi-stabilisers’ that do commute with one another.
@@ 3.1 Constructing a subsystem code from a defective stabiliser code
Figure 4: The commutation relations of a CSS subsystem code (symplectic basis). The left and right ‘columns’ are each binary matrices, each of width $2n$ (twice the number of physical qubits) and height $n$, so that, taken together, the $2n$ rows form a complete basis for $F_{2}^{2n}$. Each row of either column represents a Pauli operator. Within each row the first $n$ bits represent the places where a Pauli operator contains an $X$, and the last $n$ bits represent where it contains a $Z$. Then the diagram can be read as a series of left-right pairs of Pauli operators, so that each operator anticommutes with its neighbour and commutes with all others. Within this representation, there are blocks corresponding to the parity check matrices $H^{X}$ and $H^{Z}$, the gauge operators $G^{X}$ and $G^{Z}$, and logical operators $L^{X}$ and $L^{Z}$. To complete the basis, I’ve included the destabiliser matrices $D^{X}$ and $D^{Z}$, although for practical purposes we do not need to compute these. The bulk of the stabilisers are the existing untouched stabilisers $U^{X}$ and $U^{Z}$, with the remaining stabilisers and gauge operators formed as products of the ‘touched’ quasi-stabilisers. Viewed this way, it is clear that a subsystem code can be seen as a stabiliser code with an arbitrary partition of the logical qubits into both logical qubits and gauge qubits. The other definition of a subsystem code, via a gauge group, is illustrated with a grey box. (Note - this representation ignores phases, which we would need to include if we were to write a full stabiliser tableau—see [21])
@@ 3.2 Producing contraction schedules and finding scheduling constraints
Contraction schedules can be uniquely specified by 1) a rooted-spanning-tree subgraph of the local graph, and 2) a ‘timing’: for each edge $e$, a timestep $1\leq t_{e}\leq t$, such that no two edges that share a node are contracted at the same time, and contractions over edges more distant from the root occur before those closer to the root. In the case of $X$-type quasi-stabilisers, this means that the first CNOTs to occur are targeted on leaf nodes of the tree, and remove the support of the Pauli string on the leaf-node qubits. Subsequent CNOTs then recede the support further up the tree until it is localised at the root. Note that some rooted spanning trees have no valid timings. The spanning trees can be enumerated in linear time using a standard algorithm [28]. The timings can also be enumerated in linear time using a simple recursive algorithm (see the source).
@@ 3.3 Constructing an integer linear programming problem
ACID solves the problem using CP-SAT [12], which is known to handle large-scale scheduling problems well. The solver is given the task of trying to minimise the number of layers required to measure all the stabilisers at least once over a round. As a secondary heuristic optimisation, it tries to maximise the number of times each stabiliser is measured. The output of the solver gives us $L$, and a ‘global schedule’—a recipe for which quasi-stabilisers are to be measured in which layers and using which contraction schedule.
@@ 3.4 Assigning detectors
Stim [11], the Clifford simulator used here for the memory experiments, defines ‘detectors’ as sets of measurements whose overall parity is deterministic when no errors occur: violations of these expected parities are reported to the decoder. It is the task of the decoder to infer from the detector results which errors occurred in the circuit.
It is most helpful to think of these detectors in terms of Pauli flows [7]. For our purposes, Pauli flows are Pauli strings that propagate through our circuit in the usual way, where we can add to the support of the Pauli string (‘seed the flow’) and where we can reduce the support of the Pauli strings (‘terminate the flow’) on measurements in the correct basis. Any Pauli flow that is both seeded and terminated within the circuit can be used to define a valid detector.
@@ 4 The surface code
Figure 6: A LUCI-like diagram showing a syndrome extraction circuit of the brickwork surface code in the presence of two dropped qubits and two dropped couplers consisting of three contraction layers. I’ve followed the notation of [8], where stabilisers that are contacted in a given layer are indicated with a black depiction of the rooted spanning tree used to contract them. Dropped qubits and couplers are indicated in red. Click here to open the circuit in Shatter.
The state of the art for handling dropouts in the surface code is the LUCI framework [8], which was the original inspiration for this work. LUCI offers an intuitive way of measuring all quasi-stabilisers over $L=4$ contraction layers within a round of syndrome extraction. The technique presented in the original paper relied on the flexibility in contraction schedules offered by a (degree-4) square-grid connectivity to deal with dropped couplers and resorted to forming product stabilisers to deal with dropped qubits, with certain pathological patterns of dropped couplers treated as dropped qubits. The more recent work [13] allows LUCI to be applied to a degree-3 hexagonal/brickwork-style layout by assigning product stabilisers as soon as couplers are dropped (as well as some scheduling improvements). In both works, couplers can often be dropped without reducing the code-capacity distance $d$ of the code. In particular, in the square-grid case the loss of an isolated coupler can always be accommodated without losing distance, and in the hex-grid case we can tolerate the loss of about a third of the couplers without losing distance, with the distance reducing by one for the other couplers.